# NASP: Network Slice as a Service Platform for 5G Networks

### v2

@@ 7 Experimental Evaluation
This section presents a comprehensive experimental evaluation of the proposed NASP solution. The results are organized along two dimensions: (i) the design and deployment of E2E NSIs and (ii) their performance and lifecycle management.
@@
Table 3: Description of Edge, Metropolitan, and Central tier instances (amazonStartedWith).
@@
Overall, these experimental results validate the capability of the NASP architecture to efficiently support network‑slice design, deployment, and dynamic lifecycle management. Beyond demonstrating flexibility across four 5G scenarios, the evaluation indicates a balanced trade‑off between deployment speed, run‑time performance, and operational cost. In this context, NASP provide a viable platform for next‑generation network slicing.
@@ 8 Conclusion
Beyond the architectural contributions, the seamless integration of standardized GSMA templates, the unification of definitions across major mobile network entities to realize an E2E architecture, and the capability to instantiate network slices for Non-3GPP (non3GPP) applications open several avenues for future work. Planned efforts include: (a) integrating a physical RAN testbed to eliminate access-layer abstractions, (b) scaling the orchestration logic to support many concurrent slices, (c) embedding security and privacy assurance loops, (d) validating interoperability with commercial MANO stacks, such as ONAP and OSM, while aligning interfaces with emerging Open Radio Access Network (O-RAN) specifications in multi-operator environments, and (e) enhancing mobility management for seamless session continuity, particularly in light of the observed slice-establishment times.

### v3

@@ 7 Experimental Evaluation
This section presents a comprehensive experimental evaluation of the proposed NASP solution. The results are organized along two dimensions: (i) the design and deployment of E2E NSIs and (ii) their performance and lifecycle management. In addition to the within-testbed measurements reported in this section, we include a structured cross-study quantitative contextualization. This analysis connects the Related Work discussion with the NASP experimental results. The cited solutions differ substantially in architectural scope (e.g., CN-focused versus E2E RAN/TN/CN orchestration), lifecycle coverage, and experimental setup (e.g., simulation, emulation, real, or hybrid testbeds). Therefore, we discuss representative reported metrics from prior work and preserve their original definitions and units. Furthermore, we qualify each comparison according to its degree of alignment with NASP (Direct, Partial, Contextual, or Not comparable). This approach provides quantitative grounding without forcing equivalence when metric semantics, control boundaries, or measurement scope differ.
@@ 7.1 Slice Provisioning and Instantiation Delay
To further contextualize the provisioning results, we relate NASP’s E2E slice deployment time (22–53 s, Fig. 9) and stepwise orchestration breakdown (Fig. 10) to timings from prior studies with different scopes. The sub-200 ms instantiation reported by Scotece and others (2023) refers to localized cloud-native CNF/core startup. This evaluation serves only as a contextual lower-bound reference, not a direct E2E comparator. In contrast, Dalgitsis and others (2024) reports strategy-dependent slice deployment times of 27.0, 19.4, 21.9, $\approx 4$, and $\approx 4$ s in a federated multi-AD cloud-native 5G setup (post-federation deployment under different NF sharing/isolation strategies), which provides a partial comparator for NASP’s 22–53 s E2E provisioning. Quantitatively, these results place both studies in the seconds-to-tens-of-seconds range when orchestration scope extends beyond isolated function startup, and show that deployment time varies materially with workflow structure and sharing/isolation choices. We use the 33–38 ms (London) and 45 ms (Ireland) cloud-region latency values reported by Dalgitsis and others (2024) only as environmental context for edge-cloud experiments, not as direct slice-KPI outcomes, to avoid conflating network-path conditions with NASP’s service-level latency measurements.
7.2 Lifecycle Management and Dynamic Reconfiguration
@@
We contextualize the runtime behavior observed in NASP against representative cloud-native slicing and core-network studies, while preserving the original metric semantics of those works. Zhao et al. (2025) reports $64.2\%$ total operating-cost reduction and $45.5\%$ normalized-performance improvement, as well as converged orchestration outcomes of $4.3/1.79$ (cost / normalized performance) versus $12/1.23$ for Atlas, and dynamic adaptation after slice departure with $25.6\%$ cost reduction and $12.8\%$ performance improvement in 1 slot. These results provide a partial/contextual reference for NASP’s runtime adaptation, cost-efficiency, and service-quality-oriented evaluation dimensions. Zhao et al. (2025) also reports up to $53.9\%$ degradation (normalized performance) for comparison systems beyond 3 slices (up to 5 slices), which we use as a partial scalability reference for multi-slice concurrency effects, although the slice abstractions and resource models differ from NASP Zhao et al. (2025).
Likewise, CoreKube (Larrea et al. (2023)) reports control-plane scalability thresholds where NextEPC saturates beyond $\approx 12$ requests/s, Open5GS saturates after $\approx 120$ requests/s, and CoreKube scales beyond; a burst-handling result of 560 msgs/s from 250 UEs; resilience experiments with a 200 s critical-error interval (about $\sim 20\times$ more frequent than the referenced baseline); and AMF CPU imbalance of $91.4/1.4/1.4\%$ across three AMF pods in the K8s-Open5GS case. These results provide contextual evidence for CN/control-plane scalability and robustness, and quantitatively demonstrate that replication alone does not ensure balanced scaling without adequate traffic distribution and orchestration. Finally, we explicitly note that the disruption interval discussed in the comparison setup corresponds to the baseline environment without NASP and is not attributed to NASP-managed AMF redeployment. This part of the discussion presents a contextual comparison of reconfiguration/failure models and automation behavior rather than a direct outage benchmark.
@@ 7.3 Resource Utilization and Cost Analysis
Table 3: Description of Edge, Metropolitan, and Central tier instances (Amazon, 2024).
@@
These experimental results validate the capability of the NASP architecture to efficiently support network‑slice design, deployment, and dynamic lifecycle management. Beyond demonstrating flexibility across four 5G scenarios, the evaluation indicates a balanced trade‑off between deployment speed, run‑time performance, and operational cost. Note that our current evaluation environment uses the my5G-RANTester for RAN emulation, with a focus on end-to-end orchestration workflows. This is because NASP operates at the orchestration layer, managing the lifecycle at the network-function level rather than at the individual-flow or physical-link level. Mobility and handover procedures are executed within the RAN execution layer, whereas NASP interfaces with RAN components at the management and orchestration layer. While mobility and large-scale attachment effects influence runtime service performance and QoS metrics, they do not affect the architectural validity of the orchestration mechanisms evaluated in this work. In this context, NASP provides a viable platform for next‑generation network slicing.
The experiments consider up to 250 concurrent UEs per slice, and our experimental topology cannot fully capture topology-driven effects such as distributed domain coordination, transport-domain expansion, and end-to-end operational delays. This is because the evaluation focuses on the scalability and efficiency of the control plane and the orchestration lifecycle (e.g., slice instantiation, configuration, and end-to-end lifecycle management) in a controlled proof-of-concept environment, rather than on stressing data-plane capacity typical of large-scale mMTC deployments. In our framework, the orchestration overhead is primarily determined by the number of managed slices, instantiated network functions, and lifecycle events, rather than directly by the number of attached devices or the raw number of switches or physical links, although the scale and heterogeneity of the topology may still influence the practical orchestration latency. Therefore, the selected experimental scale enables us to capture the relevant behavior of the orchestration mechanisms as slice management complexity increases. While mMTC scenarios typically involve much larger device populations, the goal of this study is to evaluate the orchestration framework’s ability to efficiently deploy, manage, and coordinate multiple network slices. Future work will extend the evaluation to larger device populations and heterogeneous, geographically distributed topologies, including real-world RAN environments and mobility scenarios, to further assess the system under large-scale mMTC conditions.
@@ 8 Conclusion
Beyond the architectural contributions, the seamless integration of standardized GSMA templates, the unification of definitions across major mobile network entities to realize an E2E architecture, and the capability to instantiate network slices for Non-3GPP (non3GPP) applications open several avenues for future work. Planned efforts include: (a) integrating real-world physical RAN deployments to assess radio-level effects and mobility procedures that were abstracted in this study, (b) scaling the physical topology across heterogeneous domains to stress-test the data plane under massive machine-type communications (mMTC) limitations, (c) embedding advanced AI-driven analytics for predictive scaling and anomaly classification within the closed-loop assurance mechanisms, and (d) validating interoperability with commercial MANO stacks, such as ONAP and OSM, while aligning interfaces with emerging Open Radio Access Network (O-RAN) specifications in multi-operator environments.